# Hypformer: Exploring Efficient Hyperbolic Transformer Fully in Hyperbolic Space

### v1

@@ 3.2. Self-Attention Module
(4) $\mathbf{Att}_{i}\odot^{\kappa}\mathbf{V}_{j}:=$ $\frac{\sum_{j=1}^{{N}}\alpha_{ij}\mathbf{V}_{j}}{\sqrt{|{\kappa}|\left|\|\sum_{k=1}^{{N}}\alpha_{ik}\mathbf{V}_{k}\|_{\mathcal{L}}\right|}}.$
@@ 4.1. Hyperbolic Transformation with Curvatures (HTC)
(9) $\operatorname{HTC}(\mathbf{x};f_{t},\mathbf{W},\kappa_{1},\kappa_{2}):=\left(\underbrace{\sqrt{\frac{\kappa_{1}}{\kappa_{2}}\|f_{t}(\mathbf{x};\mathbf{W})\|_{2}^{2}-1/\kappa_{2}}}_{\text{time-like dimension }},\underbrace{\sqrt{\frac{\kappa_{1}}{\kappa_{2}}}f_{t}(\mathbf{x};\mathbf{W})}_{\text{space-like dimension }}\right)^{T},$
@@ 4.2. Hyperbolic Readjustment and Refinement with Curvatures (HRC)
(13) $\mathrm{HRC}(\mathbf{x};f_{r},\kappa_{1},\kappa_{2}):=\left(\underbrace{\sqrt{\frac{\kappa_{1}}{\kappa_{2}}\|f_{r}(\mathbf{x}_{[1:]})\|_{2}^{2}-1/\kappa_{2}}}_{\text{time-like dimension}},\underbrace{\sqrt{\frac{\kappa_{1}}{\kappa_{2}}}f_{r}(\mathbf{x}_{[1:]})}_{\text{space-like dimension}}\right)^{T}.$
@@ 4.3. Hyperbolic Linear Attention
[…]
@@
[…]
@@ 5.1. Experiments on Large Graphs
Experimental Findings. Table 1 summarizes the results of our experiments. Hypformer consistently outperforms other models across various large-scale graph datasets, demonstrating substantial improvements. It is worth noting that models, such as GraphFormer (Ying et al., 2021), GraphTrans (Wu et al., 2021), and GraphGPS (Rampášek et al., 2022), HAN (Gulcehre et al., 2019), HNN++ (Shimizu et al., 2020) and F-HNN (Chen et al., 2022a), have difficulty operating effectively on large-scale graph data. In addition, our method significantly outperforms the recent approaches such as, SGFormer and NodeFormer accross all tested scenarios, highlighting its superior effectivness. Importantly, Hypformer exhibits robust scalability, maintaining its performance advantage even on the largest dataset, ogbn-papers100M, where previous Transformer-based models have encountered limitations.
@@ 5.2. Experiments on Small/Medium Graphs
Table 3. Experimental results on semi-supervised classification on Mini-ImageNet and 20News-Groups where we use k-NN (with different k’s) for artificially constructing an input graph. The best and second-best results are highlighted in red bold and underlined, respectively, where models with and without graph are compared separately.
@@
LDS (Franceschi et al., 2019) OOM OOM OOM OOM ${66.15}\pm 0.36$ $64.70\pm 1.07$ $63.51\pm 0.64$ $63.51\pm 1.75$
NodeFormer (Wu et al., 2022) $\underline{86.77\pm 0.45}$ $\underline{86.74\pm 0.23}$ $\underline{86.87\pm 0.41}$ ${86.64\pm 0.42}$ ${66.01\pm 1.18}$ ${65.21\pm 1.14}$ ${64.69\pm 1.31}$ ${64.55\pm 0.97}$

### v2

# Hypformer: Exploring Efficient Transformer Fully in Hyperbolic Space

@@ 3.2. Self-Attention Module
(4) $\mathbf{Att}_{i}\odot^{\kappa}\mathbf{V}_{j}$= $\frac{\sum_{j=1}^{{N}}\alpha_{ij}\mathbf{V}_{j}}{\sqrt{|{\kappa}|\left|\|\sum_{k=1}^{{N}}\alpha_{ik}\mathbf{V}_{k}\|_{\mathcal{L}}\right|}}.$
@@ 4.1. Hyperbolic Transformation with Curvatures (HTC)
(9) $\operatorname{HTC}(\mathbf{x};f_{t},\mathbf{W},\kappa_{1},\kappa_{2})=\left(\underbrace{\sqrt{\frac{\kappa_{1}}{\kappa_{2}}\|f_{t}(\mathbf{x};\mathbf{W})\|_{2}^{2}-1/\kappa_{2}}}_{\text{time-like dimension }},\underbrace{\sqrt{\frac{\kappa_{1}}{\kappa_{2}}}f_{t}(\mathbf{x};\mathbf{W})}_{\text{space-like dimension }}\right)^{T},$
@@ 4.2. Hyperbolic Readjustment and Refinement with Curvatures (HRC)
(13) $\mathrm{HRC}(\mathbf{x};f_{r},\kappa_{1},\kappa_{2})=\left(\underbrace{\sqrt{\frac{\kappa_{1}}{\kappa_{2}}\|f_{r}(\mathbf{x}_{[1:]})\|_{2}^{2}-1/\kappa_{2}}}_{\text{time-like dimension}},\underbrace{\sqrt{\frac{\kappa_{1}}{\kappa_{2}}}f_{r}(\mathbf{x}_{[1:]})}_{\text{space-like dimension}}\right)^{T}.$
@@ 4.3. Hyperbolic Linear Attention
where $\mathcal{Q},\mathcal{K}\text{ and }\mathcal{V}\in\mathbb{L}^{N\times d^{\prime},\kappa_{2}}$ Given that the subsequent pairwise similarity computation and aggregation essentially constitute a weighted sum, and their calculation does not involve transformations on the time-like dimension, we adopted the idea of HRC to achieve this. Specifically, we first slice the values of the space-like dimension,
@@
where $\mathbf{e}\in\mathbb{L}^{d^{\prime},\kappa_{2}}$ represents the transpose of row in $\mathcal{Q},\mathcal{K}\text{ and }\mathcal{V}$. In this case, $t$ represents a scaling factor, which we set as a trainable parameter in the experiments. The focused strategy is inspired by the work in (Han et al., 2023). A $p>1$ sharpens the paired points, i.e., it enhances the similarity within each group while diminishing the similarity between the groups. Conversely, a $p<1$ has the opposite effect.
@@ 5.1. Experiments on Large Graphs
Experimental Findings. Table 1 summarizes the results of our experiments. Hypformer consistently outperforms other models across various large-scale graph datasets, demonstrating substantial improvements. It is worth noting that models, such as GraphFormer (Ying et al., 2021), GraphTrans (Wu et al., 2021), and GraphGPS (Rampášek et al., 2022), HAN (Gulcehre et al., 2019), HNN++ (Shimizu et al., 2020) and F-HNN (Chen et al., 2022a), have difficulty operating effectively on large-scale graph data. In addition, our method significantly outperforms the recent approaches such as, SGFormer and NodeFormer across all tested scenarios, highlighting its superior effectiveness. Importantly, Hypformer exhibits robust scalability, maintaining its performance advantage even on the largest dataset, ogbn-papers100M, where previous Transformer-based models have encountered limitations.
@@ 5.2. Experiments on Small/Medium Graphs
Table 3. Experimental results on semi-supervised classification on Mini-ImageNet and 20News-Groups where we use k-NN (with different k’s) for artificially constructing an input graph. The best and second-best results are highlighted in red bold and underlined, respectively.
@@
LDS (Franceschi et al., 2019) OOM OOM OOM OOM $\textbf{66.15}\pm 0.36$ $64.70\pm 1.07$ $63.51\pm 0.64$ $63.51\pm 1.75$
NodeFormer (Wu et al., 2022) $\underline{86.77\pm 0.45}$ $\underline{86.74\pm 0.23}$ $\underline{86.87\pm 0.41}$ $86.64\pm 0.42$} $\underline{66.01\pm 1.18}$ $65.21\pm 1.14$} $64.69\pm 1.31$} $64.55\pm 0.97$}